Question 1 Evaluate the integral. S22*4*3 dx • 115 + 4x375/4.c 0 - 4(5 + 4x3)-3/4.c © 2/5 + 4x3)5/4 + c 5 + 4x3) 5/4.0
Answer #1

Let I= ( 22² 4 5+423 dhe = 2/ 22 (5+423) 4 the *I=2f (5 + 42 37 4 22 de put 5+ 4223 = t 4 Bx2) de = dt >2th = dt I= 2 st 4